OpenStudy (anonymous):

Factor (x^2-2x)+1-9z^2

OpenStudy (anonymous):

does it help to know that \[x^2-2x+1=(x-1)^2\]?

OpenStudy (anonymous):

that gives you \[(x-1)^2-9z^2\] which is the difference of two squares, and you can factor as \[a^2-b^2=(a+b)(a-b)\]with \[a=(x-1)\] and \[b=3z\]

OpenStudy (anonymous):

this is the answer (x-1+3z)(x-1-3z) I need to know the steps of how to get this answer

OpenStudy (anonymous):

did you understand the last one i wrote? \[a^2-b^2=(a+b)(a-b)\] so \[(x-1)^2-9z^2=(x-1+3z)(x-1-3z)\]

OpenStudy (anonymous):

the gimmick is to recognize that \[x^2-2x+1=(x-1)^2\]

OpenStudy (anonymous):

\[\large \color{blue}a^2-\color{red}b^2=(\color{red}a+\color{blue}b)(\color{red}a-\color{blue}b)\]\[\large \color{red}{(x-1)}^2-\color{blue}{3z}^2\] \[\huge =(\color{red}{x-1}+\color{blue}{3z})(\color{red}{x-1}-\color{blue}{3z})\]

OpenStudy (anonymous):

how do factor a^3+125 using the difference of two cubes

OpenStudy (anonymous):

it is not the difference of two cubes, it is the sum of two cubes